# Topic-based Watermarks for LLM-Generated Text

### v1

@@ 4.2. Proposed Watermarking Scheme
As discussed in Section 2, LLMs generate the next token in the output text sequence from the previously generated token, specifically the work by (Kirchenbauer et al., 2023), the tokens are taken from the specified $G$ list. Our proposed watermarking scheme is inspired by this process, however our methodology differs with the utilization of extracted topics from a non-watermarked LLM-generated output and the generation of $G$ and $R$ list pairs for each topic influencing a watermarked LLM.
We propose the concept of topic-based $G$ and $R$ list pairs combined with a topic-based watermarking mechanism (TBWM). The proposed TBWM framework consists of two different LLM components, a watermarked and non-watermarked LLM. The TBWM extracts the topic from the input text sequence or a non-watermarked LLM-generated output based on the input text sequence. Figure 1 describes the proposed topic-based watermarking scheme.
@@ 4.2.1. Non-Watermarked LLM Topic Extraction
A proposed method for topic extraction employs an unwatermarked LLM. This process initiates with a user prompting the LLM with an input text sequence. Subsequently, this input sequence is introduced to the unwatermarked LLM which produces a non-watermarked output text sequence. This output serves as the bases for extracting topics relevant to the original input provided by the user. Furthermore, the non-watermarked output is then fed into a watermarked LLM in combination with a command to rephrase the text, such as “rephrase this text.” Both the non-watermarked output and the adjustments in token weight, derived from the extracted topics, are employed within the watermarked LLM to generate a watermark on the output text sequence. The utilization of the non-watermarked over the input text sequence is to extract accurate topics when (i) a small input prompt occurs, consisting of a couple words to a sentence in length and (ii) when misspellings and other grammar issues are present. The output of the non-watermarked LLM allows for a sufficient amount of quality text to be generated in order to extract an accurate topic, identifying the topic of interest of the corresponding user. Furthermore, for a single input text sequence, the watermarking scheme can extract multiple topics. For instance, in an input asking for medical injuries in different types of sports, the non-watermarked output could summarize the generalized topics from this specific output as “sports” and “medical injuries.” However, as to not degrade model performance or topic extraction quality, given a large non-watermarked output (e.g., paragraph(s)), the topic extraction may iterate over all sentences in the text sequence to analyze for relevant topics and skip those that are not strongly related to specified topics. The extracted topic or combination of topics are used for seed and $G$ and $R$ list pair generation based on the number of extracted topics. The $G$ and $R$ list pairs are utilized to tune the token weights for the watermarked LLM along with the prior non-watermarked LLM-generated output that was used for topic extraction. The watermarked LLM utilizes the non-watermarked output text sequence in combination with a rephrase command to correctly sample from the generated $G$ list per topic. Similar to the watermarking algorithm proposed by Kirchenbauer et al. (Kirchenbauer et al., 2023), we sample tokens from the $G$ list to generate an output sequence from the original input prompt.
@@
4.2.2. Input Prompt Topic Extraction
There exist two different methods for the extraction of topics: either directly from the user’s input text sequence or from an output produced by a non-watermarked LLM, which itself is based on the input text sequence. The extraction of topics from the input text is feasible when the text is free from spelling and grammatical errors and contains enough text to accurately extract topics. Simultaneously, the input text is processed through a non-watermarked LLM, resulting in an unwatermarked output based on the said input. The non-watermarked output serves as the input for the watermarked LLM along with the output and the adjustments in token weight, derived from the topics extracted from the input text sequence, are utilized within the watermarked LLM to produce a watermarked output text sequence. The approach where topics are employed to generate the $G$ and $R$ list pairs remains constant across both the non-watermarked topic extraction and the input prompt topic extraction methods.
@@ 4.3. Watermark Detection Mechanism
Existing watermarking algorithms that rely on the $G$ and $R$ list pair methodology are based on the occurrence of $G$ and $R$ listed words within the target text (LLM output text sequence) described in Section 3.2. This detection method focuses on the frequency of $R$ list violations within the target text. Text produced by human individuals typically exhibit a higher usage of $R$ listed tokens, resulting in fewer instances of $G$ listed tokens in the target text. Watermarked LLM-generated text includes significantly more $G$ listed tokens over $R$ listed tokens. Kirchenbauer et al. (Kirchenbauer et al., 2023) demonstrated the low likelihood of an individual being able to generate a text sequence, aligning with $G$ listed tokens similar to that of a watermarked LLM-generated output.
The proposed topic-based detection framework seen in Figure 2, mirrors the approach of leveraging $G$ and $R$ token list pairs. However, it introduces a novel mechanism classifying text as either human- or LLM-generated by extracting and analyzing the topics present in the input text sequence. The overall process involves the creation of specific $G$ and $R$ list pairs for each identified topic allowing the detection mechanism to iterate through all list pairs per topic to perform classification. Specifically, upon receiving an input text sequence, the watermarking framework identifies one or multiple topics present within the input text, utilizing topic modeling techniques (e.g., Latent Dirichlet allocation, non-negative factorization (Nath, 2023)). For each of the topics identified, the framework accesses the predefined pair of $G$ and $R$ lists, quantifying the occurrences of token from the pair of lists within the text. Given the distributions of $G$ and $R$ tokens across multiple different topics within the input text sequence, we employ hypothesis testing to determine the target text’s classification. The specific statistical tests chosen will depend on the distribution characteristics of the token occurrence, aimed at evaluating the confidence level, whether the target text is human- or LLM-generated. Kirchenbauer et al. (Kirchenbauer et al., 2023) utilized a “one portion z-test” involving the $G$ list of tokens with the expected value of tokens $T$ as $T/2$ and the variance of $T/4$. The overall z-test is denoted as:
@@
Figure 2. Detection of topic-based watermarking scheme. Token list pairs, generated from the extracted topics of the input text, are compared to the output text sequence of the watermarked LLM to determine the classification of the target text as human- or LLM-generated.

### v2

@@ 4.2. Proposed Watermarking Scheme
As discussed in Section 2, LLMs generate the next token in the output text sequence from the previously generated token, specifically the work by (Kirchenbauer et al., 2023), the tokens are taken from the specified $G$ list. Our proposed watermarking scheme is inspired by this process, however our methodology differs with the utilization of extracted topics from a non-watermarked LLM-generated output (or the input prompt) and the generation of $G$ and $R$ list pairs based on the extracted topic so that the detection algorithm does not need to exhaustively try all previously generated $G$ and $R$ list pairs.
We propose the concept of topic-based $G$ and $R$ list pairs combined with a topic-based watermarking mechanism (TBWM). The proposed TBWM framework consists of two different LLM components, a watermarked and non-watermarked LLM. The TBWM extracts the topic from the input prompt or a non-watermarked LLM-generated output based on the input text sequence. Figure 1 describes the proposed topic-based watermarking scheme.
@@ 4.2.1. Non-Watermarked LLM Topic Extraction
The proposed method for topic extraction employs an non-watermarked LLM output. This process initiates with a user prompting the LLM with an input prompt. Subsequently, this input sequence is input to the LLM which produces a non-watermarked output text sequence. This output serves as the basis for extracting topics relevant to the input prompt provided by the user. Furthermore, the non-watermarked output is then fed into the LLM in combination with a command to rephrase the text, such as “rephrase this text.” Both the non-watermarked output and the adjustments in token weight, derived from the extracted topics, are employed within the watermarked LLM to generate a watermark on the output text sequence. A secret key is used at this stage to influence the creation of the $G$ and $R$ list pairs. The lists are generated based on a pseudorandom function that incorporates the secret key, ensuring the watermarking scheme remains unpredictable and more robust against unauthorized detection. The utilization of the non-watermarked output over the input text sequence is to extract accurate topics when (i) a small input prompt occurs, consisting of a couple words to a sentence in length, and (ii) when misspellings and other grammar issues are present. The non-watermarked output of the LLM allows for a sufficient amount of quality text to be generated in order to extract an accurate topic, identifying the topic of interest of the corresponding user. Furthermore, for a single input text sequence, the watermarking scheme can extract multiple topics. For instance, in an input asking for medical injuries in different types of sports, the non-watermarked output could summarize the generalized topics from this specific output as “sports” and “medical injuries.” However, as to not degrade model performance or topic extraction quality, given a large non-watermarked output (e.g., paragraph(s)), the topic extraction may iterate over all sentences in the text sequence to analyze for relevant topics and skip those that are not strongly related to specified topics. The extracted topic or combination of topics are then used for seed and $G$ and $R$ list pair generation based on the number of extracted topics. The $G$ and $R$ list pairs are utilized to tune the token weights for the watermarked LLM along with the prior non-watermarked LLM-generated output that was used for topic extraction. The watermarked LLM utilizes the non-watermarked output text sequence in combination with a rephrase command to correctly sample from the generated $G$ list per topic. Similar to the watermarking algorithm proposed by Kirchenbauer et al. (Kirchenbauer et al., 2023), we sample tokens from the $G$ list to generate an output sequence from the original input prompt.
Via a similar process, we consider an alternative topic extraction method, directly from the user’s input text sequence. The extraction of topics from the input text is feasible when the text is free from spelling and grammatical errors and contains enough text to accurately extract topics. Simultaneously, the input text is processed through a non-watermarked LLM, resulting in a non-watermarked output based on the said input. The non-watermarked output serves as the input for the watermarked LLM along with the output and the adjustments in token weight, derived from the topics extracted from the input text sequence, are utilized within the watermarked LLM to produce a watermarked output text sequence. The approach where topics are employed to generate the $G$ and $R$ list pairs remains constant across both the non-watermarked topic extraction and the input prompt topic extraction methods.
@@ 4.3. Watermark Detection Mechanism
Existing watermarking algorithms that rely on the $G$ and $R$ list pair methodology are based on the occurrence of $G$ and $R$ listed words within the target text (LLM output text sequence) described in Section 3.1. This detection method focuses on the frequency of $R$ list violations within the target text. Text produced by human individuals typically exhibit a higher usage of $R$ listed tokens, resulting in fewer instances of $G$ listed tokens in the target text. Conversely, watermarked LLM-generated text includes significantly more $G$ listed tokens over $R$ listed tokens. However, it is important to mention the infeasibility of this approach due to the detection method’s need to exhaustively try different list pairs or seeds of list pairs, complicating its practical application. Kirchenbauer et al. (Kirchenbauer et al., 2023) demonstrated the low likelihood of an individual being able to generate a text sequence, aligning with $G$ listed tokens similar to that of a watermarked LLM-generated output.
The proposed topic-based detection framework shown in Figure 2, mirrors the approach of leveraging $G$ and $R$ token list pairs. However, it introduces a novel mechanism classifying text as either human- or LLM-generated by extracting and analyzing the topics present in the non-watermarked LLM output sequence. The overall process involves the creation of specific $G$ and $R$ list pairs for each identified topic allowing the detection mechanism to iterate through all list pairs per topic to perform classification. Specifically, upon receiving a non-watermarked LLM output sequence, the detection framework identifies one or multiple topics present within the input text, utilizing topic modeling techniques (e.g., Latent Dirichlet allocation, non-negative factorization) (Nath, 2023). For each of the topics identified, the framework generates the pair of $G$ and $R$ lists, quantifying the occurrences of token from the pair of lists within the text. Given the distributions of $G$ and $R$ tokens across multiple different topics within the input text sequence, we employ hypothesis testing to determine the target text’s classification. The specific statistical tests chosen depends on the distribution characteristics of the token occurrence, aimed at evaluating the confidence level, whether the target text is human- or LLM-generated. Kirchenbauer et al. (Kirchenbauer et al., 2023) utilizes a “one portion z-test” involving the $G$ list of tokens with the expected value of tokens $T$ as $T/2$ and the variance of $T/4$. The overall z-test is denoted as:
@@
Figure 2. Detection mechanism of topic-based watermarking scheme. Token list pairs, generated from the extracted topics of the non-watermarked LLM output, are compared to the output text sequence of the watermarked LLM to determine the classification of the target text as human- or LLM-generated.